# Bounds on Dao numbers and applications to regular local rings

### v1

@@ Abstract.
The so-called Dao numbers are a sort of measure of the asymptotic behaviour of full properties of certain product ideals in a Noetherian local ring $R$ with infinite residue field and positive depth. In this paper, we answer a question of H. Dao on how to bound such numbers. The auxiliary tools range from Castelnuovo-Mumford regularity of appropriate graded structures to reduction numbers of the maximal ideal. In particular, we substantially improve previous results (and answer questions) by the authors. As an application, we provide new characterizations of when $R$ is regular; for instance, we show that this holds if and only if the maximal ideal of $R$ can be generated by a $d$-sequence (in the sense of Huneke) if and only if the third Dao number of any (minimal) reduction of the maximal ideal vanishes.
@@ 1. Motivation: Dao’s problem on the fullness of certain ideals
It is worth observing that, if $(R,\mathfrak{m},K)$ and $I$ are as above, then as shown in [17, Proposition 2.2] the basic relations among the Dao numbers of $I$ are
@@
In the case where $I$ is a reduction of $\mathfrak{m}$, a lower bound for ${\mathfrak{d}}_{3}(I)$ is given by ${\rm r}_{I}(\mathfrak{m})$, the reduction number of $\mathfrak{m}$ with respect to $I$, which follows immediately from [17, Theorem 3.4] (the question as to whether ${\rm r}_{I}(\mathfrak{m})\leq{\mathfrak{d}}_{2}(I)$ remains unanswered). Otherwise, for a general $I$ (not necessarily a reduction of $\mathfrak{m}$), a more elaborated lower bound was established in [6, Proposition 1.8].
So, in the present paper, our main goal is to answer the upper bound part of Dao’s question, by using two fundamental numerical invariants in commutative algebra: the Castelnuovo-Mumford regularity (of appropriate graded structures) and, again, the reduction number. The connection between these numbers and Dao’s problem was first exploited in [17] and later developed even further (concerning specifically the Castelnuovo-Mumford regularity) in [6]. The present work establishes, in fact, generalizations and substantial improvements of the results given in these two papers. Additionally, we shall finally derive, as an application of some of our results, new characterizations of regular local rings.
@@ 2.2. Castelnuovo-Mumford regularity
We recall below a few basic rules about this invariant. For details, we refer to [2, p. 277, (a), (c) and (d)], [5, Corollary 20.19] and [10, Lemma 3.1].
@@ 2.4. Ratliff-Rush operation
This enables us to consider the following helpful number (inspired by [19, Proposition 4.2]).
@@
Let us invoke a couple of useful properties. First, according to [14, Lemma 2.2] we can write
@@
Moreover, if $\mathrm{gr}_{I}(R)=\bigoplus_{k\geq 0}I^{k}/I^{k+1}$ denotes the associated graded ring of $I$, then by [19, Remark 1.6] we get that $\widetilde{I^{k}}=I^{k}$ for all $k\geq 0$ (i.e., $s(I)=1$) if and only if
@@ 3. Upper bounds on Dao numbers via Castelnuovo-Mumford regularity
Before establishing the main results of the section, we fix a piece of notation. For a graded $R$-module $M=\bigoplus_{k\geq 0}M_{k}$ and an integer $\ell\geq 0$, we can consider the truncation $M_{\geq\ell}=\bigoplus_{k\geq\ell}M_{k}$. Specifically, with notation as in (2.5) and in case $(R,\mathfrak{m},K)$ is a local ring or a standard graded $K$-algebra, we will be interested in the truncation
@@
Here we are interested in tackling the upper bound part of Question 1.2 in terms of the Castelnuovo-Mumford regularity of appropriate graded structures. The first result in this direction, in case $R$ is local and $I$ is a reduction of $\mathfrak{m}$, was proved in [17] and can be stated as follows.
@@
(((([17, Theorem 3.10])))) Let $(R,\mathfrak{m})$ be a local ring with infinite residue field and ${\rm depth}\,R>0$, and let $I$ be a reduction of $\mathfrak{m}$. Then,
@@
Here, we prove the following result, which establishes [6, Conjecture 0.1] in its full generality.
Theorem 3.3.
Let $R$ and $I$ be as in Theorem 3.2. Then,
Proof. We may suppose $\mathfrak{d}_{3}(I)>0$, otherwise there is nothing to prove. Let $\mathfrak{M}=\mathcal{R}(\mathfrak{m})_{+}=\bigoplus_{k>0}\mathfrak{m}^{k}$ be the homogeneous maximal ideal of $\mathcal{R}(\mathfrak{m})$. It was shown in [6, proof of Theorem 1.3] that
[…]
Now, since
$\frac{(I\mathfrak{m}^{k+1}:\mathfrak{m})\cap\mathfrak{m}^{k}}{I\mathfrak{m}^{k}}\ \subset\ \frac{I\mathfrak{m}^{k+1}:\mathfrak{m}}{I\mathfrak{m}^{k}}\quad\mbox{for\, all}\quad k\geq 0,$
we have
(2) $\mathrm{Soc}(\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I))\ \subset\ \mathfrak{D}_{\mathfrak{m}}(I)$
as a graded $\mathcal{R}(\mathfrak{m})$-submodule of $\mathfrak{D}_{\mathfrak{m}}(I)$. Next, let $t=\max\{k\mid\mathfrak{D}_{\mathfrak{m}}(I)_{k}\neq 0\}$. Then, $t=\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\mathfrak{D}_{\mathfrak{m}}(I)=\mathfrak{d}_{3}(I)-1$. Because $\mathfrak{M}\cdot\mathfrak{D}_{\mathfrak{m}}(I)_{t}=0$, it follows that
$\mathfrak{D}_{\mathfrak{m}}(I)_{t}\ \subset\ [0:_{\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I)}\mathfrak{M}]_{t}\ =\ \mathrm{Soc}(\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I))_{t}$
and note that the opposite inclusion always holds by equation (2). Therefore,
$\mathrm{Soc}(\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I))_{t}=\mathfrak{D}_{\mathfrak{m}}(I)_{t}\neq 0$
and consequently
$\mathfrak{d}_{3}(I)-1\ =\ t$ $\leq\ \max\{j\mid\mathrm{Soc}(\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I))_{j}\neq 0\}$
$\leq\ \textrm{reg}_{\mathcal{R}(\mathfrak{m})}\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I)$
$=\ \textrm{reg}_{\mathcal{R}(\mathfrak{m})}\mathcal{R}(\mathfrak{m},I)-1.$
The assertion follows. $\square$
Naturally, we can ask when the inclusion (2) is actually an equality.
Proposition 3.4.
Let $R$ and $I$ be as in Theorem 3.2. If $R$ is a standard graded $K$-algebra or $\mathrm{depth}\,\mathrm{gr}_{\mathfrak{m}}(R)>0$, then $\mathrm{Soc}(\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I))=\mathfrak{D}_{\mathfrak{m}}(I)$.
Proof. In view of equation (1), it suffices to show that
$(x_{1}+J)f\in I\mathfrak{m}^{k+1}\subset\mathfrak{m}^{k+1}$
and so $\deg f+1\geq k+1$. Thus $f\in\mathfrak{m}^{k}$ and consequently $I\mathfrak{m}^{k+1}:\mathfrak{m}\subset\mathfrak{m}^{k}$ for all $k\geq 0$. Otherwise, if $\mathrm{depth}\,\mathrm{gr}_{\mathfrak{m}}(R)>0$ then, as we know, $\widetilde{\mathfrak{m}^{k}}=\mathfrak{m}^{k}$ for all $k\geq 0$. Consequently,
$\square$
We conclude this section with some other consequences of Theorem 3.3.
Corollary 3.5.
Let $R$ and $I$ be as in Theorem 3.2. If $\mathfrak{d}_{3}(I)>0$, then $\mathfrak{d}_{3}(I)$ is equal to the highest socle degree of $\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I)$ increased by one, that is,
$\mathfrak{d}_{3}(I)\ =\ \max\{j\mid\mathrm{Soc}(\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I))_{j}\neq 0\}+1.$
Proof. Set $h=\max\{j\mid\mathrm{Soc}(\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I))_{j}\neq 0\}$. In the proof of Theorem 3.3 we have seen that $\mathfrak{d}_{3}(I)\leq h+1$. On the other hand, since $\mathfrak{d}_{3}(I)>0$, we can write $\mathfrak{d}_{3}(I)=\max\{j\mid\mathfrak{D}_{\mathfrak{m}}(I)_{j}\neq 0\}+1$. By equation (2) we have an inclusion $\mathrm{Soc}(\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I))\subset\mathfrak{D}_{\mathfrak{m}}(I)$, so that
$\mathfrak{d}_{3}(I)\ =\ \max\{j\mid\mathfrak{D}_{\mathfrak{m}}(I)_{j}\neq 0\}+1\ \geq\ h+1,$
thus completing the proof. $\square$
Next, we estimate the Castelnuovo-Mumford regularity of some natural blowup algebras that are related to the Dao module of $I$.
Definition 3.6.
[…]
Notation 3.7.
For simplicity, we set
[…]
The lemma below is a special case of [25, Corollary 3].
Lemma 3.8.
Now we prove:
Corollary 3.9.
Let $R$ and $I$ be as in Theorem 3.2. Then,
In particular,
Now, suppose $\mathfrak{d}_{3}(I)>0$. In particular $\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\mathfrak{D}_{\mathfrak{m}}(I)=\mathfrak{d}_{3}(I)-1$. On the other hand, Theorem 3.3 yields
and thus the above exact sequence implies
$=\ \mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\mathcal{R}(\mathfrak{m},I).$
Finally, if we suppose that $\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\mathcal{R}(\mathfrak{m},I)>\mathfrak{d}_{3}(I)>0$, we derive
Next, we strengthen [6, Proposition 1.6] by finding the relation between the regularity of an ideal as in (2.5) and that of the ambient Rees algebra $\mathcal{R}(J)$.
@@
Notice that, as a direct consequence of Theorem 3.3 and Lemma 3.10, we rediscover Theorem 3.1.
@@
Let $(R,\mathfrak{m})$ be a local ring with infinite residue field and ${\rm depth}\,R>0$. By [20, Proposition 1.5], we have $\mathfrak{m}^{n+1}:\mathfrak{m}=\mathfrak{m}^{n}$ for all $n\geq\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\,\mathcal{R}(\mathfrak{m})$. Now set $I={\mathfrak{m}}^{k}$ for any given $k\geq 2$ (note $I$ cannot be a reduction of $\mathfrak{m}$). We can write
@@
Let $R$ be as in [17, Example 4.3]. As observed there, $\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\,\mathcal{R}(\mathfrak{m})=8$. Consider the ideal $I={\mathfrak{m}}^{2}$. By Example 3.13 above, we can write

### v2

@@ Abstract.
The so-called Dao numbers are a sort of measure of the asymptotic behaviour of full properties of certain product ideals in a Noetherian local ring $R$ with infinite residue field and positive depth. In this paper, we answer a question of H. Dao on how to bound such numbers. The auxiliary tools range from Castelnuovo-Mumford regularity of appropriate graded structures to reduction numbers of the maximal ideal. In particular, we substantially improve previous results (and answer questions) by the authors. Finally, as an application of the theory of Dao numbers, we provide new characterizations of when $R$ is regular; for instance, we show that this holds if and only if the maximal ideal of $R$ can be generated by a $d$-sequence (in the sense of Huneke) if and only if the third Dao number of any (minimal) reduction of the maximal ideal vanishes.
@@ 1. Motivation: Dao’s problem on the fullness of certain ideals
It is worth observing that, if $(R,\mathfrak{m},K)$ and $I$ are as above, then as shown in [18, Proposition 2.2] the basic relations among the Dao numbers of $I$ are
@@
In the case where $I$ is a reduction of $\mathfrak{m}$, a lower bound for ${\mathfrak{d}}_{3}(I)$ is given by ${\rm r}_{I}(\mathfrak{m})$, the reduction number of $\mathfrak{m}$ with respect to $I$, which follows immediately from [18, Theorem 3.4] (the question as to whether ${\rm r}_{I}(\mathfrak{m})\leq{\mathfrak{d}}_{2}(I)$ remains unanswered). Otherwise, for a general $I$ (not necessarily a reduction of $\mathfrak{m}$), a more elaborated lower bound was established in [6, Proposition 1.8].
So, in the present paper, our main goal is to answer the upper bound part of Dao’s question, by using two fundamental numerical invariants in commutative algebra: the Castelnuovo-Mumford regularity (of appropriate graded structures) and, again, the reduction number. The connection between these numbers and Dao’s problem was first exploited in [18] and later developed even further (concerning specifically the Castelnuovo-Mumford regularity) in [6]. The present work establishes, in fact, generalizations and substantial improvements of the results given in these two papers. Additionally, we shall finally derive new characterizations of regular local rings.
@@ 2.2. Castelnuovo-Mumford regularity
We recall below a few basic rules about this invariant. For details, we refer to [2, p. 277, (a), (c) and (d)], [5, Corollary 20.19] and [11, Lemma 3.1].
@@ 2.4. Ratliff-Rush operation
This enables us to consider the following helpful number (inspired by [20, Proposition 4.2]).
@@
Let us invoke a couple of useful properties. First, according to [15, Lemma 2.2] we can write
@@
Moreover, if $\mathrm{gr}_{I}(R)=\bigoplus_{k\geq 0}I^{k}/I^{k+1}$ denotes the associated graded ring of $I$, then by [20, Remark 1.6] we get that $\widetilde{I^{k}}=I^{k}$ for all $k\geq 0$ (i.e., $s(I)=1$) if and only if
@@ 3. Upper bounds on Dao numbers via Castelnuovo-Mumford regularity
Before establishing the results of this section, we fix a piece of notation. For a graded $R$-module $M=\bigoplus_{k\geq 0}M_{k}$ and an integer $\ell\geq 0$, we can consider the truncation $M_{\geq\ell}=\bigoplus_{k\geq\ell}M_{k}$. Specifically, using Notation 2.5 and assuming $(R,\mathfrak{m},K)$ is a local ring or a standard graded $K$-algebra, we will be interested in the truncation
@@
Here we are focused on tackling the upper bound part of Question 1.2 in terms of the Castelnuovo-Mumford regularity of appropriate graded structures. The first result in this direction, in case $R$ is local and $I$ is a reduction of $\mathfrak{m}$, was proved in [18] and can be stated as follows.
@@
(((([18, Theorem 3.10])))) Let $(R,\mathfrak{m},K)$ be a local ring with infinite residue field and ${\rm depth}\,R>0$, and let $I$ be a reduction of $\mathfrak{m}$. Then,
@@
We present the following auxiliary basic lemma, which additionally refines [6, Proposition 1.6] by establishing the relationship between the regularity of an extended ideal as in Notation 2.5 and that of the ambient Rees algebra $\mathcal{R}(J)$.
@@
Our main result establishes [6, Conjecture 0.1] in its full generality, as follows.
Theorem 3.4.
Let $R$ and $I$ be as in Theorem 3.2. Then,
Proof. Clearly, we may suppose $\mathfrak{d}_{3}(I)>0$. Set $t=\mathfrak{d}_{3}(I)-1$, and let us assume first that $s(\mathfrak{m})\leq t$. Then,
$I\mathfrak{m}^{t+1}:\mathfrak{m}\subseteq\mathfrak{m}^{t+1}:\mathfrak{m}=\widetilde{\mathfrak{m}^{t+1}}:\mathfrak{m}=\widetilde{\mathfrak{m}^{t}}=\mathfrak{m}^{t}.$
Moreover, there is a well-known isomorphism (see, for instance, [10, p. 268])
$H_{n}(\mathcal{R}(\mathfrak{m})_{+};\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I))\cong e_{1}\wedge\cdots\wedge e_{n}\,(0:_{\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I)}\mathcal{R}(\mathfrak{m})_{+}),$
where the first module is the $n$th Koszul homology module of $\mathcal{R}(\mathfrak{m})_{+}$ with respect to $\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I)$, the integer $n$ is the minimal number of generators of the irrelevant ideal $\mathcal{R}(\mathfrak{m})_{+}$, and each $e_{i}$ has degree one. Notice that we can write
(1) $0:_{\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I)}\mathcal{R}(\mathfrak{m})_{+}\ =\ \bigoplus_{k\geq 0}\frac{(I\mathfrak{m}^{k+1}:\mathfrak{m})\cap\mathfrak{m}^{k}}{I\mathfrak{m}^{k}},$
and furthermore
(2) $(0:_{\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I)}\mathcal{R}(\mathfrak{m})_{+})_{t}\ =\ \frac{(I\mathfrak{m}^{t+1}:\mathfrak{m})\cap\mathfrak{m}^{t}}{I\mathfrak{m}^{t}}\ =\ \frac{I\mathfrak{m}^{t+1}:\mathfrak{m}}{I\mathfrak{m}^{t}}\ =\ \mathfrak{D}_{\mathfrak{m}}(I)_{t}\neq 0.$
Consequently, using [2, Theorem 8.1.3] we obtain
$\mathfrak{d}_{3}(I)-1\ =\ t$ $\leq\ \max\{j\mid(0:_{\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I)}\mathcal{R}(\mathfrak{m})_{+})_{j}\neq 0\}$
$=\ \max\{j-n\mid H_{n}(\mathcal{R}(\mathfrak{m})_{+};\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I))_{j}\neq 0\}$
$\leq\ \max\{j-i\mid H_{i}(\mathcal{R}(\mathfrak{m})_{+};\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I))_{j}\neq 0\}$
$=\ \textrm{reg}_{\mathcal{R}(\mathfrak{m})}\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I)$
$=\ \textrm{reg}_{\mathcal{R}(\mathfrak{m})}\mathcal{R}(\mathfrak{m},I)-1.$
Next, assume that $t<s(\mathfrak{m})$. Since $\mathfrak{m}$ contains a regular element, we deduce from [21, Proposition 2.1(ii)] that $s(\mathfrak{m})\leq\mathrm{max}\{1,\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\,\mathcal{R}(\mathfrak{m})\}$. If $\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\,\mathcal{R}(\mathfrak{m})=0$ then $\mathrm{depth}\,\mathrm{gr}_{\mathfrak{m}}(R)>0$ (see Remark 2.10), and the result follows from [6, Theorem 1.3]. On the other hand, if $\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\,\mathcal{R}(\mathfrak{m})\geq 1$, given that $t<s(\mathfrak{m})$, we conclude
${\mathfrak{d}}_{3}(I)\leq s(\mathfrak{m})\leq\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\,\mathcal{R}(\mathfrak{m})\leq\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\,\mathcal{R}(\mathfrak{m},I),$
where the last inequality is guaranteed by Lemma 3.3. $\square$
@@
Notice that we can rediscover Theorem 3.1 as a direct consequence of Theorem 3.4 and Lemma 3.3.
@@
Note that the inclusion
$0:_{\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I)}\mathcal{R}(\mathfrak{m})_{+}\ =\ \bigoplus_{k\geq 0}\frac{(I\mathfrak{m}^{k+1}:\mathfrak{m})\cap\mathfrak{m}^{k}}{I\mathfrak{m}^{k}}\ \subseteq\bigoplus_{k\geq 0}\frac{I\mathfrak{m}^{k+1}:\mathfrak{m}}{I\mathfrak{m}^{k}}\ =\ \mathfrak{D}_{\mathfrak{m}}(I)$
always holds, and naturally we can ask when it is actually an equality.
Proposition 3.7.
Let $R$ and $I$ be as in Theorem 3.2. If $R$ is a standard graded $K$-algebra or $\mathrm{depth}\,\mathrm{gr}_{\mathfrak{m}}(R)>0$, then $0:_{\mathcal{R}(\mathfrak{m})/\mathcal{R}(\mathfrak{m},I)}\mathcal{R}(\mathfrak{m})_{+}=\mathfrak{D}_{\mathfrak{m}}(I)$.
Proof. In view of equation (1), it suffices to show that
$(x_{1}+J)f\in I\mathfrak{m}^{k+1}\subset\mathfrak{m}^{k+1}$
and so $\deg f+1\geq k+1$. Thus $f\in\mathfrak{m}^{k}$ and consequently $I\mathfrak{m}^{k+1}:\mathfrak{m}\subset\mathfrak{m}^{k}$ for all $k\geq 0$. Finally, if $\mathrm{depth}\,\mathrm{gr}_{\mathfrak{m}}(R)>0$ then, as we know, $\widetilde{\mathfrak{m}^{k}}=\mathfrak{m}^{k}$ for all $k\geq 0$. It follows that
$\square$
Next, we consider some natural blowup algebras that are related to the Dao module of $I$ and whose Castelnuovo-Mumford regularity will be compared to that of the ideal $\mathcal{R}(\mathfrak{m},I)\subset\mathcal{R}(\mathfrak{m})$.
Definition 3.8.
[…]
Notation 3.9.
For simplicity, set
[…]
The lemma below is a special case of [26, Corollary 3].
Lemma 3.10.
Now we obtain the following lower bound for $\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\mathcal{R}(\mathfrak{m},I)$.
Corollary 3.11.
Let $R$ and $I$ be as in Theorem 3.2. Then,
$\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\mathcal{R}(\mathfrak{m},I)\geq{\rm max}\{\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\mathcal{P}_{\mathfrak{m}}(I),\,\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\mathcal{Q}_{\mathfrak{m}}(I)\}\ -1.$
In particular,
Note that if $\mathfrak{d}_{3}(I)>0$ then in particular $\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\mathfrak{D}_{\mathfrak{m}}(I)=\mathfrak{d}_{3}(I)-1$. On the other hand, Theorem 3.4 yields
and thus the above exact sequence implies
$=\ \mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\mathcal{R}(\mathfrak{m},I).$
So, if now we suppose that $\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\mathcal{R}(\mathfrak{m},I)>\mathfrak{d}_{3}(I)>0$, we finally derive
@@
Let $(R,\mathfrak{m})$ be a local ring with infinite residue field and ${\rm depth}\,R>0$. By [21, Proposition 1.5], we have $\mathfrak{m}^{n+1}:\mathfrak{m}=\mathfrak{m}^{n}$ for all $n\geq\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\,\mathcal{R}(\mathfrak{m})$. Now set $I={\mathfrak{m}}^{k}$ for any given $k\geq 2$ (note $I$ cannot be a reduction of $\mathfrak{m}$). We can write
@@
Let $R$ be as in [18, Example 4.3]. As observed there, $\mathrm{reg}_{\mathcal{R}(\mathfrak{m})}\,\mathcal{R}(\mathfrak{m})=8$. Consider the ideal $I={\mathfrak{m}}^{2}$. By Example 3.12 above, we can write